# Stability and convergence of multi-product expansion splitting methods with negative weights for semilinear parabolic equations

### v1

@@ 3 MPE splitting methods for general evolutionary problems
Using the formal calculus of Lie derivatives introduced in Section 2, a $k$th-order SPE splitting operator can also be expressed as
$\mathcal{S}^{[k]}(t)=\prod_{j=1}^{m}\mathrm{e}^{a_{m+1-j}tD_{A}}\mathrm{e}^{b_{m+1-j}tD_{B}}\approx\mathcal{E}_{A+B}(t)=\mathrm{e}^{tD_{A+B}},\quad 0\leq t\leq T,$ (3.1)
@@
$\mathcal{S}^{[k]}(t)=\sum_{i=1}^{M}c_{i}\prod_{j=1}^{m_{i}}\mathrm{e}^{a_{i,m_{i}+1-j}tD_{A}}\mathrm{e}^{b_{i,m_{i}+1-j}tD_{B}}\approx\mathcal{E}_{A+B}(t)=\mathrm{e}^{tD_{A+B}},\quad 0\leq t\leq T,$ (3.2)
where the product is defined downward as (2.4) and $\mathcal{I}$ is the identity operator. Alternatively, the SPE splitting operators (3.1) and the MPE splitting operators (3.2) can be written as (1.4) and (1.5), respectively, in the form of the composition of evolution operators.
@@
$\mathcal{D}(t)=\mathcal{S}^{[k]}(t)-\mathcal{E}_{A+B}(t)=\sum_{i=1}^{M}c_{i}\prod_{j=1}^{m_{i}}\mathrm{e}^{a_{i,m_{i}+1-j}tD_{A}}\mathrm{e}^{b_{i,m_{i}+1-j}tD_{B}}-\mathrm{e}^{tD_{A+B}},\quad 0<t\leq T,$ (3.3)
@@
To analyze the local splitting error of SPE, the BCH formula is an efficient tool. Regardless of whether $A$ and $B$ are linear or nonlinear, the corresponding Lie operators $D_{A}$ and $D_{B}$ are linear by definition. In fact, for any pair of linear operators $X,Y$, the BCH formula provides a series expansion for the product of exponentials:
$\mathrm{e}^{X}\mathrm{e}^{Y}=\mathrm{e}^{X+Y+\frac{1}{2}[X,Y]+\frac{1}{12}([X,X,Y]+[Y,Y,X])+\frac{1}{24}[X,Y,Y,X]+\cdots}.$ (3.5)
Here, we adopt the notation of the commutator $[X,Y]:=XY-YX$, with higher-order commutators recursively defined as $[X,X,Y]=[X,[X,Y]]$. For example, the Lie–Trotter splitting operators are given by
$\mathcal{S}^{[1],1}(t)=\mathrm{e}^{tD_{B}}\mathrm{e}^{tD_{A}},\quad\mathcal{S}^{[1],2}(t)=\mathrm{e}^{tD_{A}}\mathrm{e}^{tD_{B}}.$ (3.6)
By BCH formula and Taylor expansion (2.13), the defect operator is
$\mathcal{D}^{[1],1}(t)=\mathcal{S}^{[1],1}(t)-\mathcal{E}_{A+B}(t)=\frac{1}{2}t^{2}[D_{A},D_{B}]+O(t^{3}).$ (3.7)
Analogously, the Strang splitting operator $\mathcal{S}_{2}(t)$ is expressed as
$\mathcal{S}^{[2],1}(t)=\mathrm{e}^{\frac{t}{2}D_{A}}\mathrm{e}^{tD_{B}}\mathrm{e}^{\frac{t}{2}D_{A}}\quad\text{or}\quad\mathcal{S}^{[2],2}(t)=\mathrm{e}^{\frac{t}{2}D_{B}}\mathrm{e}^{tD_{A}}\mathrm{e}^{\frac{t}{2}D_{B}}.$ (3.8)
In a similar way,
$\mathcal{D}^{[2],1}(t)=\mathcal{S}^{[2],1}(t)-\mathcal{E}_{A+B}(t)$ $=\mathrm{e}^{tE_{1}+t^{3}E_{3}+t^{5}E_{5}+\cdots}-\mathrm{e}^{t(D_{A}+D_{B})}$ (3.9)
$=t^{3}E_{3}+\frac{1}{2}t^{4}(E_{1}E_{3}+E_{3}E_{1})+O(t^{5}),$
where $E_{i}$ are the nested commutators of $D_{A}$ and $D_{B}$ as follows [50]:
$E_{1}=$ $D_{A}+D_{B},\quad E_{3}=\frac{1}{12}[D_{B},D_{B},D_{A}]-\frac{1}{24}[D_{A},D_{A},D_{B}],$ (3.10)
$E_{5}=$ $\frac{7}{5760}[D_{A},D_{A},D_{A},D_{A},D_{B}]-\frac{1}{720}[D_{B},D_{B},D_{B},D_{B},D_{A}]+\frac{1}{360}[D_{A},D_{B},D_{B},D_{B},D_{A}]$
$+\frac{1}{360}[D_{B},D_{A},D_{A},D_{A},D_{B}]-\frac{1}{480}[D_{A},D_{A},D_{B},D_{B},D_{A}]+\frac{1}{120}[D_{B},D_{B},D_{A},D_{A},D_{B}].$
Since $D_{A}$ and $D_{B}$ generally do not commute, i.e., $D_{A}D_{B}\neq D_{B}D_{A}$, the local splitting error is unavoidable.
To derive arbitrarily high-order conditions, an effective algorithm for the BCH formula and related calculations in the Lyndon basis are proposed in [11] for SPE splittings. Although BCH formula can effectively reveal error terms for SPE splittings, it becomes excessively cumbersome when applied to MPE splitting methods, particularly for higher-order terms. Therefore, a more practical approach is to directly use the Taylor expansion (2.13) of defect operator (3.3) to verify and derive higher-order splitting methods.
3.1 Direct construction from order conditions
Some interesting MPE splitting methods have been proposed and studied in literature. For example, two third-order splitting operators were proposed in [10]:
@@ 3.1 Direct construction from order conditions
$\mathcal{S}^{[4],1}\left(t\right)=$ $\frac{2}{3}\mathrm{e}^{\frac{1}{2}tD_{B}}\mathrm{e}^{\frac{1}{2}tD_{A}}\mathrm{e}^{\frac{1}{2}tD_{B}}\mathrm{e}^{\frac{1}{2}tD_{A}}+\frac{2}{3}\mathrm{e}^{\frac{1}{2}tD_{A}}\mathrm{e}^{\frac{1}{2}tD_{B}}\mathrm{e}^{\frac{1}{2}tD_{A}}\mathrm{e}^{\frac{1}{2}tD_{B}}$ (3.13)
$-\frac{1}{6}\mathrm{e}^{tD_{A}}\mathrm{e}^{tD_{B}}-\frac{1}{6}\mathrm{e}^{tD_{B}}\mathrm{e}^{tD_{A}}.$
We now present the order conditions for constructing high-order MPE splittings, which are actually consistent with those for ordinary differential systems established in [42]. The following conditions are required to ensure that the coefficients of the identity operator and the first-order Taylor-expansion terms $D_{A},D_{B}$ vanish:
$\sum^{M}_{i=1}c_{i}=1,\quad\sum^{M}_{i=1}c_{i}\sum_{j=1}^{m_{i}}a_{i,j}=1,\quad\sum^{M}_{i=1}c_{i}\sum_{j=1}^{m_{i}}b_{i,j}=1.$ (3.14)
The following conditions are required to ensure that the coefficients of the second-order Taylor-expansion terms $D_{A}D_{A},D_{B}D_{B},D_{A}D_{B}$ and $D_{B}D_{A}$ vanish:
$\begin{gathered}\sum^{M}_{i=1}c_{i}\left(\sum_{j=1}^{m_{i}}a_{i,j}\right)^{2}=1,\quad\sum^{M}_{i=1}c_{i}\left(\sum_{j=1}^{m_{i}}b_{i,j}\right)^{2}=1,\\ \sum_{i=1}^{M}c_{i}\sum_{j_{1}=1}^{m_{i}}a_{i,j_{1}}\sum_{j_{2}=j_{1}}^{m_{i}}b_{i,j_{2}}=\frac{1}{2},\quad\sum_{i=1}^{M}c_{i}\sum_{j_{1}=1}^{m_{i}-1}b_{i,j_{1}}\sum_{j_{2}=j_{1}+1}^{m_{i}}a_{i,j_{2}}=\frac{1}{2}.\end{gathered}$ (3.15)
The following conditions are required to ensure that the coefficients of the third-order Taylor-expansion terms $D_{A}D_{A}D_{A}$, $D_{B}D_{B}D_{B}$, $D_{A}D_{A}D_{B}$, $D_{B}D_{A}D_{A}$, $D_{A}D_{B}D_{A}$, $D_{B}D_{A}D_{B}$, $D_{A}D_{B}D_{B}$, and $D_{B}D_{B}D_{A}$ vanish:
$\begin{gathered}\sum^{M}_{i=1}c_{i}\left(\sum_{j=1}^{m_{i}}a_{i,j}\right)^{3}=1,\quad\sum^{M}_{i=1}c_{i}\left(\sum_{j=1}^{m_{i}}b_{i,j}\right)^{3}=1,\\ \sum_{i=1}^{M}c_{i}\left(\sum_{j_{1}=1}^{m_{i}}a_{i,j_{1}}^{2}\sum_{j_{2}=j_{1}}^{m_{i}}b_{i,j_{2}}+2\sum_{j_{1}=1}^{m_{i}-1}a_{i,j_{1}}\sum_{j_{2}=j_{1}+1}^{m_{i}}a_{i,j_{2}}\sum_{j_{3}=j_{2}}^{m_{i}}b_{i,j_{3}}\right)=\frac{1}{3},\\ \sum_{i=1}^{M}c_{i}\left(\sum_{j_{1}=1}^{m_{i}-1}b_{i,j_{1}}\sum_{j_{2}=j_{1}+1}^{m_{i}}a_{i,j_{2}}^{2}+2\sum_{j_{1}=1}^{m_{i}-2}b_{i,j_{1}}\sum_{j_{2}=j_{3}+1}^{m_{i}-1}a_{i,j_{2}}\sum_{j_{3}=j_{2}+1}^{m_{i}}a_{i,j_{3}}\right)=\frac{1}{3},\\ \sum_{i=1}^{M}c_{i}\sum_{j_{1}=1}^{m_{i}-1}a_{i,j_{1}}\sum_{j_{2}=j_{1}}^{m_{i}-1}b_{i,j_{2}}\sum_{j_{3}=j_{2}+1}^{m_{i}}a_{i,j_{3}}=\frac{1}{6},\\ \sum_{i=1}^{M}c_{i}\sum_{j_{1}=1}^{m_{i}-1}b_{i,j_{1}}\sum_{j_{2}=j_{1}+1}^{m_{i}}a_{i,j_{2}}\sum_{j_{3}=j_{2}}^{m_{i}}b_{i,j_{3}}=\frac{1}{6},\\ \sum_{i=1}^{M}c_{i}\left(\sum_{j_{1}=1}^{m_{i}}a_{i,j_{1}}\sum_{j_{2}=j_{1}}^{m_{i}}b_{i,j_{2}}^{2}+2\sum_{j_{1}=1}^{m_{i}-1}a_{i,j_{1}}\sum_{j_{2}=j_{1}}^{m_{i}-1}b_{i,j_{2}}\sum_{j_{3}=j_{2}+1}^{m_{i}}b_{i,j_{3}}\right)=\frac{1}{3},\\ \sum_{i=1}^{M}c_{i}\left(\sum_{j_{1}=1}^{m_{i}-1}b_{i,j_{1}}^{2}\sum_{j_{2}=j_{1}+1}^{m_{i}}a_{i,j_{2}}+2\sum_{j_{1}=1}^{m_{i}-2}b_{i,j_{1}}\sum_{j_{2}=j_{1}+1}^{m_{i}-1}b_{i,j_{2}}\sum_{j_{3}=j_{2}+1}^{m_{i}}a_{i,j_{3}}\right)=\frac{1}{3}.\end{gathered}$ (3.16)
For higher-order conditions, the coefficient expressions for each Taylor-expansion term can be derived through coefficient matching. Naturally, for an operator splitting method to achieve the order of accuracy $k$, it must satisfy all conditions up to and including the order $k$. Specifically, there are $12^{k+1}-$ 1 order conditions for $k$th-order MPE splitting methods.
In order to simplify these conditions, we consider the MPE splitting operators of the following form:
$\mathcal{S}^{[k]}(t)=\sum_{i=1}^{M}c_{i}\mathcal{S}^{[k]}_{i}(t),$ (3.17)
where $\mathcal{S}^{[k]}_{i}(t)$ consists of an odd number of evolution operators and have the following symmetric form
$\mathcal{S}^{[k]}_{i}(t)=$ $\frac{1}{2}\mathrm{e}^{d_{i,1}tD_{A}}\mathrm{e}^{d_{i,2}tD_{B}}\cdots\mathrm{e}^{d_{i,2m_{i}}tD_{B}}\mathrm{e}^{d_{i,2m_{i}+1}tD_{A}}$ (3.18)
$+\frac{1}{2}\mathrm{e}^{d_{i,1}tD_{B}}\mathrm{e}^{d_{i,2}tD_{A}}\cdots\mathrm{e}^{d_{i,2m_{i}}tD_{A}}\mathrm{e}^{d_{i,2m_{i}+1}tD_{B}},$
with $d_{i,j}=d_{i,2m_{i}+2-j}\geq 0$ for $j=1,2,\cdots,2m_{i}+1$ and $\sum_{j=0}^{m_{i}}d_{i,1+2j}=\sum_{j=1}^{m_{i}}d_{i,2j}=1$.
The form (3.18) exhibits three useful properties for reducing order conditions. First, it preserves symmetry under the exchange of operators, meaning that swapping $D_{A}$ and $D_{B}$ does not alter the overall structure. Consequently, the order conditions are identical. For example, the order conditions for Taylor-expansion terms $D_{A}D_{B}D_{A}$ and $D_{B}D_{A}D_{B}$ are the same. Second, the splitting operator enforces the symmetry of the step coefficients, expressed as $d_{i,j}=d_{i,2m_{i}+2-j}$. This property ensures that if one Taylor-expansion terms is the reverse of another, their order conditions are identical. For example, Taylor-expansion terms $D_{A}D_{B}D_{B}$ and $D_{B}D_{B}D_{A}$ share the same order conditions. This structural symmetry reduces the number of distinct order conditions that must be considered. The third property that could be helpful for higher-order constructions is time reversibility. SPE splitting operators possessing this property have been studied in [50], resulting in the following lemma.
Lemma 3.1 (Time reversibility [50]).
Let $\mathcal{S}^{[k]}(t)$ be an SPE operator of the form (3.1), which possesses the time reversibility
$\mathcal{S}^{[k]}(t)\mathcal{S}^{[k]}(-t)=\mathcal{S}^{[k]}(-t)\mathcal{S}^{[k]}(t)=\mathcal{I}.$ (3.19)
Assume that $\mathcal{S}^{[k]}(t)$ is expanded as
$\mathcal{S}^{[k]}(t)=\mathrm{e}^{tE_{1}+t^{2}E_{2}+t^{3}E_{3}+t^{4}E_{4}+\cdots},$ (3.20)
where the error terms $E_{i}$ are the nested commutators of $D_{A}$ and $D_{B}$ depending on the specific form of $\mathcal{S}^{[k]}(t)$. Then, the following holds:
$E_{2}=E_{4}=E_{6}=\cdots=0.$ (3.21)
Note that both components of $\mathcal{S}^{[k]}(t)$ defined in (3.18) exhibit time reversibility. Consequently, if all components of the MPE splitting operators adhere to the structure of (3.18), some even-order conditions can be naturally disregarded.
By leveraging the aforementioned order conditions, incorporating the special form (3.18), and utilizing time reversibility, we can systematically construct the following two new fourth-order splitting operators:
@@
$\mathcal{S}^{[4],3}\left(t\right)=$ $\frac{4}{3}\mathrm{e}^{\frac{1}{8}tD_{A}}\mathrm{e}^{\frac{1}{4}tD_{B}}\mathrm{e}^{\frac{3}{8}tD_{A}}\mathrm{e}^{\frac{1}{2}tD_{B}}\mathrm{e}^{\frac{3}{8}tD_{A}}\mathrm{e}^{\frac{1}{4}tD_{B}}\mathrm{e}^{\frac{1}{8}tD_{A}}$ (3.23)
$+\frac{4}{3}\mathrm{e}^{\frac{1}{8}tD_{B}}\mathrm{e}^{\frac{1}{4}tD_{A}}\mathrm{e}^{\frac{3}{8}tD_{B}}\mathrm{e}^{\frac{1}{2}tD_{A}}\mathrm{e}^{\frac{3}{8}tD_{B}}\mathrm{e}^{\frac{1}{4}tD_{A}}\mathrm{e}^{\frac{1}{8}tD_{B}}$
@@ 3.2 Construction by Richardson extrapolation
Richardson extrapolation is another efficient and convenient approach to construct high-order MPE splitting methods. For completeness, we summarize and elaborate on this approach as presented in [16]. Since the Strang splitting operator $\mathcal{S}_{2}(t)$ in (3.8) satisfies the time reversibility (3.19), it follows from (3.1) that the defect operator $\mathcal{D}(t)$ in (3.9) contains only odd power terms of $t$. To be specific,
$\mathcal{S}_{2}(t)$ $=\exp\left[t(D_{A}+D_{B})+t^{3}E_{3}+t^{5}E_{5}+\cdots\right]$ (3.24)
$=\mathrm{e}^{t(D_{A}+D_{B})}+t^{3}E_{3}+\frac{1}{2}t^{4}\left[(D_{A}+D_{B})E_{3}+E_{3}(D_{A}+D_{B})\right]+O\left(t^{5}\right).$
The error terms $E_{i}$ are identical to those in (3.10), representing the nested commutators of $D_{A}$ and $D_{B}$. Thus, as a basic component of MPE splitting operators, one of the Strang splitting operators $\mathcal{S}_{2}(t)$ in (3.8) can be used to construct a $2k$-order approximation $\mathcal{S}^{[2k]}(t)$ via Richardson extrapolation. This implies that the $\gamma$-th power of $\mathcal{S}_{2}(t)$ with a step size of $t/\gamma$ must take the form
γ $=\exp\left[t(D_{A}+D_{B})+\gamma^{-2}t^{3}E_{3}+\gamma^{-4}t^{5}E_{5}+\cdots\right]$ (3.25)
$=\mathrm{e}^{t(D_{A}+D_{B})}+\gamma^{-2}t^{3}E_{3}+\frac{1}{2}\gamma^{-2}t^{4}\left[(D_{A}+D_{B})E_{3}+E_{3}(D_{A}+D_{B})\right]+O\left(t^{5}\right).$
Then, a Richardson extrapolation yields
$\frac{\gamma^{2}[\mathcal{S}_{2}(t/\gamma)]^{\gamma}-\mathcal{S}_{2}(t)}{\gamma^{2}-1}=\mathrm{e}^{t(D_{A}+D_{B})}+O\left(t^{5}\right),$ (3.26)
where both the third and fourth order errors of $\mathcal{S}_{2}(t/\gamma)$ can be eliminated simultaneously, yielding a fourth-order splitting operator. Similarly, the leading $12k+$ 1 and $22k+$ 2 order errors are multiplied by $\gamma^{-2k}$ and can be eliminated together. Thus, for a given set of distinct $k$ numbers $\{\gamma_{i}\}_{i=1}^{k}$, one can have a $2k$-th order approximation
$\mathrm{e}^{t(D_{A}+D_{B})}=\sum_{i=1}^{k}c_{i}\left[\mathcal{S}_{2}\left(\frac{t}{\gamma_{i}}\right)\right]^{\gamma_{i}}+O\left(t^{2k+1}\right),$ (3.27)
provided that $c_{i}$ satisfy the simple Vandermonde equation:
$\left(\begin{array}[]{ccccc}1&1&1&\ldots&1\\ \gamma_{1}^{-2}&\gamma_{2}^{-2}&\gamma_{3}^{-2}&\ldots&\gamma_{k}^{-2}\\ \gamma_{1}^{-4}&\gamma_{2}^{-4}&\gamma_{3}^{-4}&\ldots&\gamma_{k}^{-4}\\ \vdots&\vdots&\vdots&&\vdots\\ \gamma_{1}^{-2(k-1)}&\gamma_{2}^{-2(k-1)}&\gamma_{3}^{-2(k-1)}&\ldots&\gamma_{k}^{-2(k-1)}\end{array}\right)\left(\begin{array}[]{c}c_{1}\\ c_{2}\\ c_{3}\\ \vdots\\ c_{k}\end{array}\right)=\left(\begin{array}[]{c}1\\ 0\\ 0\\ \vdots\\ 0\end{array}\right)$ (3.28)
To be specific, for a given set of distinct $k$ numbers $\{\gamma_{i}\}_{i=1}^{k}$,
@@
$\mathcal{S}^{[2k]}\left(t\right):=\sum^{k}_{i=1}c_{i}\left[\mathcal{S}_{2}\left(\frac{t}{\gamma_{i}}\right)\right]^{\gamma_{i}},\quad\text{with}\quad c_{i}=\prod^{k}_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}\frac{\gamma_{i}^{2}}{\gamma_{i}^{2}-\gamma_{j}^{2}}.$ (3.29)
The natural sequence $\{\gamma_{i}:\gamma_{i}=i\}_{i=1}^{k}$ produces a $2k$-th order splitting operator (for example, $k=2,3,4,5$):
$\mathcal{S}^{[4],4}(t)$ $:=-\frac{1}{3}\mathcal{S}_{2}(t)+\frac{4}{3}\left[\mathcal{S}_{2}\left(\frac{t}{2}\right)\right]^{2},$ (3.30)
$\mathcal{S}^{[6]}(t)$ $:=\frac{1}{24}\mathcal{S}_{2}(t)-\frac{16}{15}\left[\mathcal{S}_{2}\left(\frac{t}{2}\right)\right]^{2}+\frac{81}{40}\left[\mathcal{S}_{2}\left(\frac{t}{3}\right)\right]^{3},$
$\mathcal{S}^{[8]}(t)$ $:=-\frac{1}{360}\mathcal{S}_{2}(t)+\frac{16}{45}\left[\mathcal{S}_{2}\left(\frac{t}{2}\right)\right]^{2}-\frac{729}{280}\left[\mathcal{S}_{2}\left(\frac{t}{3}\right)\right]^{3}+\frac{1024}{315}\left[\mathcal{S}_{2}\left(\frac{t}{4}\right)\right]^{4},$
$\mathcal{S}^{[10]}(t)$ $:=\frac{1}{8640}\mathcal{S}_{2}(t)-\frac{64}{945}\left[\mathcal{S}_{2}\left(\frac{t}{2}\right)\right]^{2}+\frac{6561}{4480}\left[\mathcal{S}_{2}\left(\frac{t}{3}\right)\right]^{3}$
$\quad-\frac{16384}{2835}\left[\mathcal{S}_{2}\left(\frac{t}{4}\right)\right]^{4}+\frac{390625}{72576}\left[\mathcal{S}_{2}\left(\frac{t}{5}\right)\right]^{5}.$
4 Uniform stability of MPE splitting methods for semilinear parabolic equations

### v2

@@ 3 Some MPE splitting methods
Using the formal calculus of Lie derivative introduced in (A), a $k$th-order splitting operator can also be expressed as
$\mathcal{S}^{[k]}(t)=\prod_{j=1}^{m}\mathrm{e}^{a_{m+1-j}tD_{A}}\mathrm{e}^{b_{m+1-j}tD_{B}}\approx\mathcal{E}_{A+B}(t)=\mathrm{e}^{tD_{A+B}},\qquad 0\leq t\leq T,$ (3.1)
@@
$\mathcal{S}^{[k]}(t)=\sum_{i=1}^{M}c_{i}\prod_{j=1}^{m_{i}}\mathrm{e}^{a_{i,m_{i}+1-j}tD_{A}}\mathrm{e}^{b_{i,m_{i}+1-j}tD_{B}}\approx\mathcal{E}_{A+B}(t)=\mathrm{e}^{tD_{A+B}},\qquad 0\leq t\leq T,$ (3.2)
where the product is defined downward as (2.2) and $\mathcal{I}$ is the identity operator. Alternatively, the splitting operators (3.1) and the MPE splitting operators (3.2) can be written as (1.4) and (1.5), respectively, in the form of the composition of evolution operators. It should also be noted that, different from (1.4) and (1.5), the evolution operators in (3.1) and (3.2) acts in reversed order (from right to left).
@@
$\mathcal{D}(t)=\mathcal{S}^{[k]}(t)-\mathcal{E}_{A+B}(t)=\sum_{i=1}^{M}c_{i}\prod_{j=1}^{m_{i}}\mathrm{e}^{a_{i,m_{i}+1-j}tD_{A}}\mathrm{e}^{b_{i,m_{i}+1-j}tD_{B}}-\mathrm{e}^{tD_{A+B}},\qquad 0<t\leq T,$ (3.3)
@@
Some interesting MPE splitting methods have been proposed and studied in the literature. For example, two third-order splitting operators were proposed in [12]:
@@
$\mathcal{S}^{[4],1}\left(t\right)=$ $\frac{2}{3}\mathrm{e}^{\frac{1}{2}tD_{B}}\mathrm{e}^{\frac{1}{2}tD_{A}}\mathrm{e}^{\frac{1}{2}tD_{B}}\mathrm{e}^{\frac{1}{2}tD_{A}}+\frac{2}{3}\mathrm{e}^{\frac{1}{2}tD_{A}}\mathrm{e}^{\frac{1}{2}tD_{B}}\mathrm{e}^{\frac{1}{2}tD_{A}}\mathrm{e}^{\frac{1}{2}tD_{B}}-\frac{1}{6}\mathrm{e}^{tD_{A}}\mathrm{e}^{tD_{B}}-\frac{1}{6}\mathrm{e}^{tD_{B}}\mathrm{e}^{tD_{A}}.$ (3.7)
Using the order conditions, we can also construct the following two new fourth-order splitting operators.
@@
$\mathcal{S}^{[4],3}\left(t\right)=$ $\frac{4}{3}\mathrm{e}^{\frac{1}{8}tD_{A}}\mathrm{e}^{\frac{1}{4}tD_{B}}\mathrm{e}^{\frac{3}{8}tD_{A}}\mathrm{e}^{\frac{1}{2}tD_{B}}\mathrm{e}^{\frac{3}{8}tD_{A}}\mathrm{e}^{\frac{1}{4}tD_{B}}\mathrm{e}^{\frac{1}{8}tD_{A}}+\frac{4}{3}\mathrm{e}^{\frac{1}{8}tD_{B}}\mathrm{e}^{\frac{1}{4}tD_{A}}\mathrm{e}^{\frac{3}{8}tD_{B}}\mathrm{e}^{\frac{1}{2}tD_{A}}\mathrm{e}^{\frac{3}{8}tD_{B}}\mathrm{e}^{\frac{1}{4}tD_{A}}\mathrm{e}^{\frac{1}{8}tD_{B}}$ (3.9)
@@
Richardson extrapolation is another efficient and convenient approach to construct high-order MPE splitting methods. Using the time reversibility of Strang splitting $\mathcal{S}_{2}(t)$ in (1.7), for a given set of $k$ distinct numbers $\{\gamma_{i}\}_{i=1}^{k}$, the $2k$-th order MPE splitting can be constructed [16] as follows
@@
$\mathcal{S}^{[2k]}\left(t\right)=\sum^{k}_{i=1}c_{i}\left[\mathcal{S}_{2}\left(\frac{t}{\gamma_{i}}\right)\right]^{\gamma_{i}},\qquad\text{with}\qquad c_{i}=\prod^{k}_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}\frac{\gamma_{i}^{2}}{\gamma_{i}^{2}-\gamma_{j}^{2}},~k\geq 1.$ (3.10)
The natural sequence $\{\gamma_{i}:\gamma_{i}=i\}_{i=1}^{k}$ produces $2k$-th order splittings (for example, $k=2,3,4,5$):
$\mathcal{S}^{[4],4}(t)$ $=-\frac{1}{3}\mathcal{S}_{2}(t)+\frac{4}{3}\left[\mathcal{S}_{2}\left(\frac{t}{2}\right)\right]^{2},$ (3.11)
$\mathcal{S}^{[6]}(t)$ $=\frac{1}{24}\mathcal{S}_{2}(t)-\frac{16}{15}\left[\mathcal{S}_{2}\left(\frac{t}{2}\right)\right]^{2}+\frac{81}{40}\left[\mathcal{S}_{2}\left(\frac{t}{3}\right)\right]^{3},$
$\mathcal{S}^{[8]}(t)$ $=-\frac{1}{360}\mathcal{S}_{2}(t)+\frac{16}{45}\left[\mathcal{S}_{2}\left(\frac{t}{2}\right)\right]^{2}-\frac{729}{280}\left[\mathcal{S}_{2}\left(\frac{t}{3}\right)\right]^{3}+\frac{1024}{315}\left[\mathcal{S}_{2}\left(\frac{t}{4}\right)\right]^{4},$
$\mathcal{S}^{[10]}(t)$ $=\frac{1}{8640}\mathcal{S}_{2}(t)-\frac{64}{945}\left[\mathcal{S}_{2}\left(\frac{t}{2}\right)\right]^{2}+\frac{6561}{4480}\left[\mathcal{S}_{2}\left(\frac{t}{3}\right)\right]^{3}$
$\qquad-\frac{16384}{2835}\left[\mathcal{S}_{2}\left(\frac{t}{4}\right)\right]^{4}+\frac{390625}{72576}\left[\mathcal{S}_{2}\left(\frac{t}{5}\right)\right]^{5}.$
Note that the above mentioned MPE splittings have positive time steps but possibly negative weights.
Remark 3.1.
Some other efficient MPE splitting methods with negative time steps have recently been proposed in [11, 9]. However, we will see in the following content that for these MPE splittings, the stability analysis in this work seems not applicable.
4 Stability of MPE splitting methods for semilinear parabolic equations